Heidelberg weekly meeting

Europe/Zurich
03.410 (Physikalisches Institut, Heidelberg)

03.410

Physikalisches Institut, Heidelberg

Description

We will have the meeting in HeiCONF: https://heiconf.uni-heidelberg.de/sch-nqu-dd2   Access Code: 858824

Vidyo will be used as a back-up. Room extension 10823320 or use this link.

 

Tentative minute taker list at https://docs.google.com/spreadsheets/d/15oo_a_1WPxPBOtlH0LDEKG1128y1O2lfswpUAvPCQOY/edit#gid=0

    • 14:00 14:45
      Round Table 45m

      Reports of work progress

      Speakers: Adrian Herkert (Physikalisches Institut, Uni Heidelberg), Alena Weber (Universität Heidelberg), Andre Schoening (University Heidelberg, Institute of Physiscs), Annie Meneses Gonzalez (PhD Student), Benjamin Weinläder (PI Heidelberg), David Maximilian Immig (Universität Heidelberg), Dohun Kim (Physikalisches Institut University of Heidelberg), Frank Meier Aeschbacher (Universität Heidelberg), Heiko Augustin (PI Uni Heidelberg), Lars Noehte (PI Heidelberg), Luigi Vigani (University of Oxford), Lukas Mandok (Physikalisches Institut Heidelberg), Marius Menzel, Sebastian Dittmeier (Physikalisches Institut, Universität Heidelberg), Sebastian Preuß (PI Uni Heidelberg), Thomas Theodor Rudzki (Physikalisches Institut Heidelberg)

      Round Table

      Heiko:

      -  Bochum brings equipment to Heidelberg, picks chips here

      - Delivery will be soon

      -Goal: Setup more MuPix10 Setups

      - Cleanup of lab ongoing (count inventory)

      Alena:

      -Testing hit counting chip -> works well

      -Strontium measurement with HitPix ongoing

      -working on Documentation for HitPix

      Alex:

      -Frontend adapter card development almost done

      -Cooling of equalizer chips discussed -> First 3D Model done

      Benjamin:

      -started clean-up of the lab

      -started list of all motherboards (almost complete)

      Dohun:

      -handed in Thesis! Congratulations!

      Florian:

      -connected the Oscilloscope to the PC and can control it with C++

      -DAQ and Oscilloscope can not run simultaneously at the moment

      Annie:

      -added new counting to old map ( Ivans documentation)

      -work with version 5 of standard amplifier, compare pixel sizes

      Thomas:

      -measurement is foreseen for November

      -glued two PCBs with 6 MuPix10 chips

      - gluing pattern changed and will be tested soon

      -shopping list: (Qtips, tissues, label tape, HAMEGs,….)

      -If someone wants to add to the shopping list -> contact Thomas

      Luigi:

      -working on adapter card

      André

      -Magnet successfully ramped to 1T (well centered!)

       

    • 14:45 14:55
      AOB